Find all points on the graph of the equation x^2-y^2=2x+4y where the tangent line is horizontal. Does the graph have any vertical asymptotes?
1
Expert's answer
2020-07-09T19:53:45-0400

x2-y2=2x+4y

y2+4y=x2-2x

y2+4y+4-4=x2-2x+1-1

(y+2)2 =(x-1)2+3 conjugate hyperbole

two branches

y=((x-1)2+3)1/2-2

y=-((x-1)2+3)1/2-2

find the derivatives

y`=(x-1)/((x-1)2+3)1/2

y`=-(x-1)/((x-1)2+3)1/2

the derivative is 0 for

x=1 y1=(3)1/2-2 y2=-(3)1/2-2





The line x = x0 cannot be a vertical asymptote if the function is continuous at the point x = x0. Therefore, vertical asymptotes should be sought at the points of discontinuity of the function.